# Summarization of Opinionated Political Documents with Varied Perspectives

### v1

@@ Abstract
Global partisan hostility and polarization has increased, and this polarization is heightened around presidential elections. Models capable of generating accurate summaries of diverse perspectives can help reduce such polarization by exposing users to alternative perspectives. In this work, we introduce a novel dataset and task for independently summarizing each political perspective in a set of passages from opinionated news articles. For this task, we propose a framework for evaluating different dimensions of perspective summary performance. We benchmark 10 models of varying sizes and architectures through both automatic and human evaluation. While recent models like GPT-4o perform well on this task, we find that all models struggle to generate summaries faithful to the intended perspective. Our analysis of summaries focuses on how extraction behavior depends on the features of the input documents.
@@
Nicholas Deas, Kathleen McKeown Department of Computer Science, Columbia University
@@ 1 Introduction
People who hold political ideologies tend to develop misperceptions of groups with opposing political opinions Chambers et al. (2006), and political events, such as the 2024 US presidential election, French legislative election, or the Brexit referendum, reinforce negative attitudes Hanna et al. (2013); Wellings et al. (2024). These misperceptions, reinforced by news consumption on social media Levy (2021), contribute to increased polarization and instability Braley et al. (2023); Lees and Cikara (2021). Exposure to alternative perspectives, however, has been shown to help alleviate polarization Balietti et al. (2021). To encourage such exposure, some groups focus on aggregating (e.g., All Sides [^1] https://www.allsides.com/) or summarizing different political perspectives on divisive issues (e.g., The Flip Side  [^2] https://www.TheFlipSide.io/, Ground News [^3] https://ground.news/).
@@
To evaluate current models’ capabilities in summarizing different perspectives, we propose a novel dataset and task, PoliSum [^4] We make our dataset available to researchers who have been granted permission by The Flip Side.. Our task involves generating independent abstractive summaries of both left and right-leaning political perspectives given collections of opinionated news passages with mixed perspectives. PoliSum thus requires systems to both distinguish the opinions reflected in input texts as well as produce a pair of summaries representing those differing opinions. We summarize our primary contributions as follows:
@@
    We introduce a novel benchmark dataset, PoliSum, for independently summarizing political perspectives on controversial issues, provided passages from news and op-eds.
@@
    We present and discuss benchmark evaluation results, finding that models struggle with faithfulness to the intended perspective and that improving automatic evaluation metrics is an important area for future work.
@@
    We analyze model extraction behavior (e.g., which documents models extract from and how much) and find that generated summaries suffer from biases due to input document position, document length, and use of arousing terms.
@@ 2 PoliSum Dataset
CoCoTrip Reviews $48\cdot 3$ 132.9 22.8 72.4 91.4
PoliSum Op-Eds $907\cdot 2$ 19.2 40.4 86.8 97.5
Table 1: Summary statistics for PoliSum and existing summarization datasets, including data size, reference length, and percentage of novel n-grams in the summary compared to source documents. # Gold for multi-summary datasets is presented as # Samples $\cdot$ # References/Sample. CoCoTrip statistics sourced from Iso et al. (2022).
Data Source. To enable evaluating political perspective summaries, we collect opinionated texts and paired summaries from The Flip Side [^5] We thank The Flip Side for generously allowing us to use their archive for this work: https://www.TheFlipSide.io/archives. Since September 2017, The Flip Side has covered politically controversial issues (i.e. overturning of Roe v. Wade, presidential elections) in the United States. Toward bridging political divides, each publication includes a headline detailing the issue, curated passages authored by political commentators with varying perspectives, and a pair of short summaries of the left and right perspectives [^6] Some publications lack individual perspective summaries, primarily on topics where both parties largely agree or for meta-articles that do not cover political topics. These samples are not included in the dataset.. We collect this information from The Flip Side, gathering nearly 1,000 samples. An excerpted example of source passages and summaries is shown in Table 2.
@@
PoliSum Task. Using this dataset, the aim of the PoliSum task is to generate a pair of summaries (i.e., Summaries in Table 2) representing the left and right political perspectives on a controversial topic (i.e. Headline). As input, models are provided with two sets of opinionated editorial passages (i.e., Source Passages), each representing one of the political perspectives. Thus, the dataset enables a multi-document, multi-summary task. As with traditional news summarization tasks, the summary for each perspective should condense the information present in the passages. At the same time, the generated left and right-perspective summaries should accurately represent the perspectives reflected in their respective passages.
@@ 3 Experiments
Traditional opinion summarization typically involves summarizing a common perspective among all inputs, which we refer to as the Single-Perspective setting [^7] Experimental details and results of the Single-Perspective task on PoliSum are included in Appendix E. We instead introduce the additional challenge of identifying input documents relevant to the intended perspective, referred to as the Mixed-Perspective setting (illustrated in Figure 1). In this setting, approaches are provided with a mixed set of left and right-leaning texts as input and are tasked with producing independent left and right-perspective summaries from the same input. On average, sets of inputs contain $\sim$10 passages total, $\sim$5 passages for each perspective.
@@
3.1 Baselines
A variety of zero-shot summarization baselines and instruction-tuned LLMs are considered.
Extractive Upper Bound. To represent possible extractive approaches, we calculate an upper bound performance based on similarity to the reference summaries. We consider all sentences contained in the source passages to be potential extractive summaries. We then select the sentence with the highest BERTScore relative to the reference summary and calculate the remaining performance metrics.
Zero-Shot Baselines. As zero-shot summarization baselines, we evaluate the large variants of BART Lewis et al. (2019), T5 Raffel et al. (2019), and Flan-T5 Chung et al. (2022). As BART is not pretrained on any summarization task, we use BART finetuned on the CNN/DailyMail (CNN/DM; Hermann et al. 2015; Nallapati et al. 2016) dataset. In contrast to PoliSum, CNN/DM and other prior summarization datasets lack summaries for individual perspectives. So, in order to encourage BART to generate distinct summaries matching the references’ format, the BART decoder is forced to begin generations with "The left" or "The right" respectively. T5 and Flan-T5 are prompted to generate summaries of individual perspectives. See Appendix A for other generation hyperparameter details.
LLM Baselines. We evaluate 10 LLMs of varying sizes and architectures to establish baselines for current state-of-the-art models [^8] We use instruction-tuned variants of all applicable LLMs, but for simplicity, omit ”Instruct” in names throughout results.. Mistral-7B (mistralai/Mistral-7B-Instruct-v0.1; Jiang et al. 2023), Llama-3-8B and 70B (meta-llama/Meta-Llama-3-8B-Instruct and meta-llama/Meta-Llama-3-70B-Instruct; Dubey et al. 2024), Vicuna-7B and 13B (lmsys/vicuna-7b-v1.5 and lmsys/vicuna-13b-v1.5; Chiang et al. 2023), Mixtral-8x22B (mistralai/Mixtral-8x22B-Instruct-v0.1; Jiang et al. 2024), and GPT-4o (gpt-4o-2024-05-13) [^9] https://openai.com/index/hello-gpt-4o/. See Appendix B for additional generation hyperparameters, prompts, and reproducibility details.
@@ 3.2 Metrics
Summary coverage is first measured with Rouge-1 and Rouge-L between the generated perspective summaries and their references Lin (2004a). To account for the weaknesses of Rouge, we also employ BERTScore Zhang* et al. (2020) and Bleurt Sellam et al. (2020) using the deberta-large-xnli and Bleurt-20-D6 checkpoints respectively due to higher correlations with human judgments compared to alternatives [^10] See human correlations for BERTScore linked in https://github.com/Tiiiger/bert_score and for Bleurt in https://github.com/google-research/Bleurt/blob/master/checkpoints.md. For each metric, we also use a two-sample t-test to determine whether left and right-perspective coverage scores significantly differ ($p\leq 0.05$).
@@
We measure how well the generated summaries are consistent and faithful to the inputs. Because faithfulness metrics for our specific task do not exist, we rely on existing factuality metrics as approximations: SummaC Laban et al. (2022) and AlignScore Zha et al. (2023). SummaC is a NLI-based metric that measures consistency between summaries and source documents. Alternatively, AlignScore uses information alignment to measure factual consistency. All metrics produce a score between 0 and 100 where higher values reflect more consistent and faithful summaries.
@@
To measure how well the generated summaries reflect the intended perspective, we operationalize dimensions of how perspectives are captured in summaries. Stance represents whether the summary accurately portrays the intended perspective’s overall attitude toward the object (e.g., for or against), while Object represents whether the summary identifies the salient stance object among input documents. These dimensions correspond to the attitude and stance object of the stance triangle theory respectively Du Bois (2007). Because our task requires expressing perspectives in natural language as opposed to prior stance detection work, we include two additional dimensions: Intensity reflects whether the extent to which a perspective is held or the strength of the belief is accurately captured by a perspective summary (e.g., "praises" rather than "supports" to reflect a stronger belief); and Reasoning refers to whether the elaboration in a perspective summary is supported by the input documents. Finally, our task requires summarizing multiple opinions toward a single object, which may disagree. Therefore, we include a final dimension to measure whether perspective summaries appropriately express the extent to which input documents agree or disagree, called Consensus.
@@
Additionally, we introduce initial metrics to automate selected dimensions. To measure Stance, we employ JointCL, a zero-shot stance classifier Liang et al. (2022); Allaway and McKeown (2023). Using the headlines as the object, we calculate the Root Mean Squared Error (RMSE) between predictions on the generated summaries and references. To capture Object, we use a T5-based keyphrase generation model Li et al. (2023) as a zero-shot stance object classifier, and calculate BERTScore between predictions on generated summaries and input passages. Finally, similar to Lee et al. (2022), we measure Intensity by calculating the average arousal score of polarizing terms according to the VAD lexicon Mohammad (2018). Intensity scores are again calculated as the RMSE between arousal scores for generated summaries and input passages. Due to their complexity, we leave automated evaluations of Consensus and Reasoning to future work. See Appendix F and Appendix D for further details on human and automatic metrics respectively.
@@
Table 3: Average coverage and faithfulness scores for the Mixed-Perspective setting. Best performance on each metric is bolded and second best underlined. For each model and metric, significantly higher left than right scores are highlighted in blue, and significantly higher right than left scores are highlighted in red. Higher scores reflect better performance for all metrics.
@@ 4.1 Summarization Evaluation
Table 3 presents coverage and faithfulness metrics for all models in the mixed-perspective setting. All models outperform the extractive baseline in coverage, emphasizing that abstractive approaches are vital to performing well on this task. Among abstractive approaches, Llama-3 (8B) tends to perform best on ROUGE while GPT-4o tends to perform best on Bleurt and BERTScore. Interestingly, while the larger, 13B-parameter Vicuna model performs better than its 7B-parameter variant, the same is not true for Llama-3. This suggests that larger models do not always guarantee better mixed-perspective summarization performance.
Trends in faithfulness metrics differ from that of the coverage metrics. Vicuna models achieve the best performance on SummaC for both perspectives and AlignScore for the right-leaning perspective, whereas Llama-3 (70B) is best on AlignScore for the left perspective. Overall, however, all models score far lower on faithfulness metrics than the extractive baseline, suggesting improvements and evaluation of faithfulness on this task are important issues to address in future work.
In examining differences between perspectives, BLEURT and BERTScore are significantly higher for left-leaning summaries across most models, while SummaC and AlignScore are significantly higher for most right-leaning summaries. Notably, all four metrics are model-based. These differences suggest that in addition to political biases studied in prior work (e.g., Feng et al. 2023), model-based metrics may also suffer from biases.
@@ 4.2 Perspective Evaluation
Ref. 2.9 3.2 2.9 2.9 1.8 17.7 0.68
Mistral 2.4 3.2 2.5 2.5 1.9 41.6 0.84
Mixtral 3.0 3.4 3.0 3.0 2.6 39.9 0.72
GPT-4o 3.2 3.5 3.2 3.2 2.9 33.7 0.66
Table 5: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Best scores are bolded. Higher scores reflect better performance for all metrics. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
Automatic perspective metrics are included in Table 4, and human perspective judgments of Mistral, Mixtral, GPT-4o, and reference summaries are included in Table 5 (breakdown of judgment scores and correlations are included in Appendix F). GPT-4o summaries tend to score highest on both automatic and human scores, even exceeding reference summary judgments. While the references are similar in extractiveness to GPT-4o, the summaries are drastically shorter, perhaps making the references harder to judge. Compared to human judgments, the automatic metrics for Stance, Object, and Intensity largely rank the three models in the same order, although there is little sample-level correlation with human judgments. While the metrics may capture system-level differences (i.e., between GPT-4o and Mistral), they often fail to distinguish individual summaries of similar quality (i.e., summaries by the same model). Improving metrics for measuring similarity between perspective dimensions and exploring alternative methods for measuring perspective dimensions are vital areas for future work.
5 What Drives Model Performance?
@@
5.1 Qualitative Summary Evaluation
Examples summaries generated by Mistral, Mixtral, and GPT-4o are shown in Table 6. In these examples, models appear to attend to the correct perspective within input documents, suggesting that models may have sufficient knowledge of left and right-leaning political beliefs and rhetoric to distinguish inputs in some cases. In summarizing the perspectives, however, models tend to rely on specific information only present in some of the input documents with little abstractive synthesis of the information (6(a)). Additionally, models consistently extract or paraphrase from the same documents across many examples, such as in 6(b). Such consistency suggests that qualities of the input documents themselves may drive models’ tendency to extract from them, which we investigate in subsection 5.2.

### v2

@@ Abstract
Global partisan hostility and polarization has increased, and this polarization is heightened around presidential elections. Models capable of generating accurate summaries of diverse perspectives can help reduce such polarization by exposing users to alternative perspectives. In this work, we introduce a novel dataset and task for independently summarizing each political perspective in a set of passages from opinionated news articles. For this task, we propose a framework for evaluating different dimensions of perspective summary performance. We benchmark 11 summarization models and LLMs of varying sizes and architectures through both automatic and human evaluation. While recent models like GPT-4o perform well on this task, we find that all models struggle to generate summaries that are faithful to the intended perspective. Our analysis of summaries focuses on how extraction behavior is impacted by features of the input documents [^1] We make our code available at https://github.com/NickDeas/PoliSum.
@@
Nicholas Deas, Kathleen McKeown Department of Computer Science, Columbia University [ndeas,kathy]@cs.columbia.edu
@@ 1 Introduction
Political ideologies can lead people to develop misperceptions of groups with opposing opinions Chambers et al. (2006), and political events, such as the 2024 US presidential election, French legislative election, or the Brexit referendum, can reinforce negative attitudes Hanna et al. (2013); Wellings et al. (2024). These misperceptions, reinforced by news consumption on social media Levy (2021), contribute to increased polarization and instability Braley et al. (2023); Lees and Cikara (2021). Exposure to alternative perspectives, however, has been shown to help alleviate polarization Balietti et al. (2021). To encourage such exposure, some groups focus on aggregating (e.g., All Sides [^2] https://www.allsides.com/) or summarizing different political perspectives on divisive issues (e.g., The Flip Side [^3] https://www.TheFlipSide.io/, Ground News [^4] https://ground.news/).
@@
To evaluate current models’ capabilities in summarizing multiple perspectives in a set of input documents, we propose a novel dataset, PoliSum, and task, mixed perspective summarization [^5] We make our dataset available to researchers who have been granted permission to use the data for their work by The Flip Side.. Our task involves generating independent abstractive summaries of both left and right-leaning political perspectives given collections of opinionated news passages with mixed perspectives. Mixed perspective summarization thus requires models to both distinguish the perspectives reflected in input texts as well as produce a pair of summaries representing those differing perspectives. We summarize our primary contributions as follows:
@@
    We introduce a novel benchmark dataset, PoliSum, for mixed perspective summarization, requiring models to independently summarize political perspectives on controversial issues, provided passages from news and op-eds.
@@
    We benchmark 11 models, finding that most approaches struggle to faithfully summarize the intended perspective and that improving automatic evaluation metrics is an important area for future work.
@@
    We analyze model extraction behavior (e.g., which documents models extract from and to what degree) and find that generated summaries suffer from biases due to input document position, document length, and use of arousing terms.
@@ 2 PoliSum Dataset
CoCoTrip Reviews 48348$\cdot$ 348 ⋅ 3 132.9 22.8 72.4 91.4
PoliSum Op-Eds 9072907$\cdot$ 2907 ⋅ 2 19.7 40.4 86.8 97.5
Table 1: Summary statistics for PoliSum and existing summarization datasets, including data size, reference length, and percentage of novel n-grams in the summary compared to source documents. # Gold for multi-summary datasets is presented as # Samples $\cdot$ # References/Sample. CoCoTrip statistics referenced from Iso et al. (2022).
Data Source. To enable evaluating political perspective summaries, we collect opinionated texts and paired summaries from The Flip Side [^6] We thank The Flip Side for generously allowing us to use their archive for this work: https://www.TheFlipSide.io/archives. Since September 2017, The Flip Side has covered politically controversial issues (i.e. overturning of Roe v. Wade, presidential elections) in the United States. Toward bridging political divides, each publication includes a headline reflecting the central issue, curated passages authored by political commentators with varying political leaning, and a pair of short summaries of the left and right perspectives [^7] Some publications lack individual perspective summaries, primarily on topics where both parties largely agree or for meta-articles that do not cover political topics. These samples are not included in the dataset.. We collect this information from The Flip Side archive, gathering nearly 1,000 samples. An excerpted example of source passages and summaries is shown in Table 2.
@@
Perspective Summarization Task. Using this dataset, the aim of the our proposed task is to generate a pair of summaries (i.e., Summaries in Table 2) representing the left and right political perspectives on a controversial topic (i.e. Headline). As input, models are provided with two sets of opinionated editorial passages (i.e., Source Passages), each representing one of the two political perspectives. Thus, the dataset enables a multi-document, multi-summary task. As with traditional news summarization tasks, the summary for each perspective should condense the information present in the passages. At the same time, the generated left and right-perspective summaries should faithfully represent the perspectives reflected in their respective passages.
@@ 3 Experiments
Traditional opinion summarization typically involves summarizing a common perspective among all inputs, which we refer to as the Single-Perspective setting [^8] Experimental details and results of the Single-Perspective task on PoliSum are included in Appendix E. We instead introduce an additional challenge requiring models to identify input documents relevant to the intended perspective, referred to as the Mixed-Perspective setting and illustrated in Figure 1. In this setting, approaches are provided with a mixed set of left and right-leaning passages as input and are tasked with producing independent left and right-perspective summaries from the same input. On average, sets of inputs contain $\sim$10 passages total, $\sim$5 passages for each perspective.
@@
3.1 Models
A variety of summarization models and instruction-tuned LLMs are considered.
Extractive Upper Bound. To represent potential extractive summarization approaches, we estimate an upper bound performance based on semantic similarity to the reference summaries. We consider all sentences contained in the source passages to be potential extractive summaries. We then select the sentence with the highest BERTScore relative to the reference summary and calculate the remaining coverage and faithfulness metrics.
Neural Summarization Baselines. To represent neural summarization models, we evaluate the large variants of BART Lewis et al. (2019), T5 Raffel et al. (2019), and Flan-T5 Chung et al. (2022). As BART is not pretrained on any summarization task, we use BART finetuned on the CNN/DailyMail (CNN/DM; Hermann et al. 2015; Nallapati et al. 2016) dataset. In contrast to PoliSum, CNN/DM and other prior summarization datasets lack summaries for individual perspectives. So, in order to encourage BART to generate distinct summaries matching the references’ format, the BART decoder is forced to begin generations with "The left" or "The right" respectively. T5 and Flan-T5 are prompted to generate summaries of individual perspectives. See Appendix A for other generation hyperparameter details.
LLMs. We evaluate 8 additional LLMs of varying sizes and architectures to establish baselines for current state-of-the-art models [^9] We use instruction-tuned variants of all applicable LLMs, but for simplicity, omit ”Instruct” in names throughout results.. Mistral-7B (mistralai/Mistral-7B-Instruct-v0.1; Jiang et al. 2023), Llama-3.0-8B and 70B (meta-llama/Meta-Llama-3-8B-Instruct and meta-llama/Meta-Llama-3-70B-Instruct; Dubey et al. 2024), Llama-3.3-70B (meta-llama/Meta-Llama-3.3-70B-Instruct), Vicuna-7B and 13B (lmsys/vicuna-7b-v1.5 and lmsys/vicuna-13b-v1.5; Chiang et al. 2023), Mixtral-8x22B (mistralai/Mixtral-8x22B-Instruct-v0.1; Jiang et al. 2024), and GPT-4o (gpt-4o-2024-05-13) [^10] https://openai.com/index/hello-gpt-4o/. Due to the length of input documents, we evaluate all LLMs in a zero-shot setting. See Appendix B for additional generation hyperparameters, prompts, and reproducibility details.
@@ 3.2 Metrics
Summary coverage is first measured with Rouge-1 and Rouge-L between the generated perspective summaries and their corresponding references Lin (2004a). To account for the weaknesses of Rouge, we also employ BERTScore Zhang et al. (2020) and Bleurt Sellam et al. (2020) using the deberta-large-xnli and Bleurt-20-D6 checkpoints respectively due to higher correlations with human judgments compared to alternatives [^11] See human correlations for BERTScore linked in https://github.com/Tiiiger/bert_score and for Bleurt in https://github.com/google-research/Bleurt/blob/master/checkpoints.md. For each metric, we also use a two-sample t-test to determine whether left and right-perspective coverage scores significantly differ ($p\leq 0.05$).
@@
In addition to coverage, we also measure how well the generated summaries are consistent and faithful to the input passages representing the intended perspective. Because faithfulness metrics for our specific task do not exist, we rely on existing factuality metrics as approximations: SummaC Laban et al. (2022) and AlignScore Zha et al. (2023). SummaC is a NLI-based metric that measures consistency between summaries and source documents. Alternatively, AlignScore uses information alignment to measure factual consistency. All metrics produce a score between 0 and 100 where higher values reflect more consistent and faithful summaries.
@@
To measure how well the generated summaries reflect the intended perspective, we decompose dimensions of how perspectives are captured in summaries. Stance represents whether the summary accurately portrays the intended perspective’s overall attitude toward the object (e.g., for or against), while Object represents whether the summary identifies the salient stance object among input documents. These dimensions correspond to the attitude and stance object of the stance triangle theory respectively Du Bois (2007). Because our task requires expressing perspectives in natural language as opposed to prior stance detection work, we include two additional dimensions: Intensity reflects whether the extent to which a perspective is held or the strength of the belief is accurately captured by a perspective summary (e.g., "praises" rather than "supports" may reflect a stronger belief); and Reasoning refers to whether the elaboration in a perspective summary is supported by the input documents. Finally, our task requires summarizing multiple perspectives toward a single object, which may not fully agree. Therefore, we include a final dimension to measure whether perspective summaries appropriately express the extent to which input documents agree or disagree, called Consensus.
@@
Additionally, we propose initial metrics to automate selected dimensions drawing on prior work in stance detection and political summarization. To measure Stance, we employ JointCL, a zero-shot stance classifier Liang et al. (2022); Allaway and McKeown (2023). Using the headlines as the object, we calculate the Root Mean Squared Error (RMSE) between predictions on the generated summaries and references. To capture Object, we use a T5-based keyphrase generation model Li et al. (2023) as a zero-shot stance object classifier, and calculate BERTScore between predictions on generated summaries and input passages. Finally, similar to Lee et al. (2022), we measure Intensity by calculating the average arousal score of polarizing terms according to the VAD lexicon Mohammad (2018). Intensity scores are again calculated as the RMSE between arousal scores for generated summaries and input passages. Due to their complexity, we leave further investigation of automated evaluations for Consensus and Reasoning to future work. See Appendix F and Appendix D for further details on human and automatic metrics respectively.
@@
Table 3: Average coverage and faithfulness scores for the Mixed-Perspective setting. Best model performance on each metric is bolded and second best underlined. For each model and metric, significantly higher left than right scores are highlighted in blue, and significantly higher right than left scores are highlighted in red. Higher scores reflect better performance for all metrics.
@@ 4.1 Summarization Evaluation
Table 3 presents coverage and faithfulness metrics for all models in the mixed-perspective setting. All models outperform the extractive baseline in coverage, emphasizing that abstractive approaches are vital to performing well on this task. Among abstractive approaches, Llama-3.0 (8B) tends to perform best on ROUGE while GPT-4o tends to perform best on Bleurt and BERTScore. Interestingly, while the larger, 13B-parameter Vicuna model performs better than its 7B-parameter variant, the same is not true for Llama-3.0. This suggests that larger models do not always guarantee better mixed-perspective summarization performance.
Trends in faithfulness metrics differ from that of the coverage metrics. Vicuna models achieve the best performance on SummaC for both perspectives and AlignScore for the right-leaning perspective, whereas Llama-3.3 (70B) is best on AlignScore for the left perspective. Overall, however, all models score far lower on faithfulness metrics than the extractive baseline, suggesting improving model faithfulness to the intended perspective for this task is an important issue to address in future work.
Between perspectives, BLEURT and BERTScore are significantly higher for left-leaning summaries across most models, while SummaC and AlignScore are significantly higher for most right-leaning summaries. Notably, all four metrics are model-based as opposed to ROUGE metrics which exhibit few significant differences between left and right-leaning summaries. Because the differences are consistent across most approaches, and prior works have uncovered political biases in models similar to those underlying each metric (e.g., Feng et al. 2023), it suggests that these metrics may also suffer from political biases.
@@ 4.2 Perspective Evaluation
Ref. 2.9 3.2 2.9 2.9 1.8 19.7 0.68
Mistral 2.4 3.2 2.5 2.5 1.9 48.6 0.84
Mixtral 3.0 3.4 3.0 3.0 2.6 46.2 0.72
GPT-4o 3.2 3.5 3.2 3.2 2.9 39.7 0.66
Table 5: Average human judgments of each perspective dimension with average length and extraction coverage over all inputs. Best scores are bolded. Higher scores reflect better performance for all metrics. Length is measured in Llama-3 tokens. S: Stance, O: Object, I: Intensity, C: Consensus, R: Reasoning.
@@
Automatic perspective metrics are included in Table 4, and human perspective judgments of Mistral, Mixtral, GPT-4o, and reference summaries are included in Table 5 (breakdown of judgment scores and correlations with automatic metrics are included in Appendix F and Appendix G). GPT-4o summaries tend to score highest on both automatic and human scores, even exceeding reference summary judgments. While the references are similar in extractiveness to GPT-4o, the summaries are significantly shorter which may make them difficult to accurately compare.
Attorney General William Barr Testifies
Side Summary Align SummaC
Left The left believes that Attorney General Barr’s handling of the Mueller report was a deliberate attempt to mislead the public and protect President Trump from accountability, and that Congress must take action to hold Trump accountable for his abuses of power. 0.8 22.2
Right The right believes that Attorney General Barr acted appropriately in releasing a summary of the Mueller report, and that Democrats are unfairly attacking him for not spinning the report in their favor, while also downplaying the fact that Mueller’s report did not find collusion and did not recommend obstruction charges. 94.4 24.4
Table 7: Example pair of left and right-leaning summaries generated by Llama-3.0 (70B) with AlignScore and SummaC scores.
Compared to human judgments, the automatic metrics for Stance, Object, and Intensity largely rank the three models in the same order, although there is little sample-level correlation with human judgments. While the metrics may capture system-level differences (i.e., between GPT-4o and Mistral), they often fail to distinguish individual summaries of similar quality (i.e., summaries by the same model). Improving metrics for measuring similarity between perspective dimensions and exploring alternative methods for measuring perspective dimensions are vital areas for future work.
5 What Drives Model Performance?
5.1 Qualitative Evaluation
Examples summaries generated by Mistral, Mixtral, and GPT-4o are shown in Table 6. In these examples, models appear to attend to the correct perspective within input documents, suggesting that models may have sufficient knowledge of left and right-leaning political beliefs and rhetoric to distinguish inputs in some cases. In summarizing the perspectives, however, models consistently tend to rely primarily on specific information present in a small subset of input documents with little abstractive synthesis of the information (6(a)). Additionally, models consistently extract or paraphrase from the same documents across many examples, such as in 6(b). This consistent behavior suggests that qualities of the input documents themselves may drive models’ tendency to extract from them, which we investigate in subsection 5.2.